OpenStudy (anonymous):

3xsquared-5x+3=0 need steps pls

OpenStudy (anonymous):

15x^2-19x+6 = 0 15x^2 - 10x - 9x + 6 = 0 5x(3x - 2) - 3(3x - 2) = 0 (3x - 2) (5x - 3)=0

OpenStudy (anonymous):

That is not one of my choices

OpenStudy (anonymous):

wat are your choices?

OpenStudy (anonymous):

ok.......5+or -I square root of 11 divided by 6, -5plus or minus square root of 11 divided by 6 or 5 plus or minus square root of 11 divided by 6 or -5 plus or minus I square root of 11 divided by 6